OpenStudy (anonymous):

If f(x) = x + 4 and g(x) = 2x - 3, find (g - f)(2).

jimthompson5910 (jim_thompson5910):

(g - f)(2) is the same as g(2) - f(2)

jimthompson5910 (jim_thompson5910):

are you able to find the value of g(2) ?

OpenStudy (anonymous):

u plug in 2 for x and then subtract the functions

OpenStudy (anonymous):

f(2)= 2+4 g(2)= 2(2) -3

OpenStudy (anonymous):

f(2)= 6 g(2) = 1

OpenStudy (anonymous):

ur gonna have to subtract f from g so it's 1-6

OpenStudy (anonymous):

1-6= -5